Foot of the perpendicular from a given line.

Grade:12

1 Answers

Rakhi Mukherjee Sarkar
13 years ago

Let, A(x2,y2) is the feet of a perpendicular from the point B(x1,y1) of the straight line ax + by + c = 0 then,

(x2-x1)/a = (y2-y1)/b =-2(ax1 +by1 +c)/(a2+b2).

If image of "A" w.r.t the line mirror ax + by + c = 0 is "B" then also the above is true.